A 40.8 mol sample of methane gas CH4 is places in a 1020 L container at 298K.

What is the pressure, in atm of the methane gas?
1
Expert's answer
2017-11-22T13:10:07-0500
Use the ideal gas law. From original equation PV=nRT there is P=nRT/V,
R=8.31441J/(K∙mol)=0.08206(L∙atm)/(K∙mol), i.e. P (atm) = (40.8 mol*0.08206(L∙atm)/(K∙mol)*298K)/1020 L = 0.9782 atm

Answer: 0.9782 atm